OpenStudy (anonymous):

The functions f(x) and g(x) are described below: f(x) = 2x + 9 g(x) = 2x - 15 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units.

OpenStudy (anonymous):

I think these are both parallel lines because the slope is 2

OpenStudy (bee_see):



OpenStudy (mathmate):

In general, if q(x)=f(x)+k, then q(x) is higher than f(x) by k units. (watch sign of k). If k is negative, then q(x) is lower.

OpenStudy (anonymous):

@Bee_see why solve for n?

OpenStudy (anonymous):

Sorry mathmate I don't understand what your saying.

OpenStudy (anonymous):

I thought it would be go down 8, but my dad thinks its 24

OpenStudy (bee_see):

we are going to go down from 9, right? so n is how many times you go down to get to -15. you can also do a number like if you want

OpenStudy (anonymous):

Ohh, so the answer is 24

OpenStudy (bee_see):

yes

OpenStudy (anonymous):

okay thanks :D

OpenStudy (mathmate):

New position - old position = shift = g(x)-f(x) = -15-(9)=-24 , negative means down, so down 24 units.